OpenStudy (anonymous):

i made a pizza for school (not real) and its diameter is 20 its radius is 10 the arc length is 3.92 it is 16 slices of pizza i need to find the interior angle of one slice how do i find the interior angle?

OpenStudy (anonymous):

and i need the formula for finding the area of one slice not the answer the formula

OpenStudy (anonymous):

The arc length is extraneous information, that is, it is not needed to solve this problem. You only need divide the total number of degrees (360) that are in a circle by the number of slices of pizza in order to get the interior angle for one slice.

OpenStudy (anonymous):

ok thank you what about finding the area of one slice?

OpenStudy (anonymous):

For that, you do need some of the other information. Area of circle (whole pizza) = (pi)(r^2) Area of one slice = (pi)(r^2)/16

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

You're welcome. Good luck and thx for the recognition!